### v3

# LLM one-shot style transfer for Authorship Attribution and Verification

Abstract
Computational stylometry studies writing style through quantitative textual patterns, enabling applications such as authorship attribution, identity linking, and plagiarism detection. Existing supervised and contrastive approaches often rely on datasets with spurious correlations, conflating style with topic. Despite the relevance of language modeling to these tasks, the pre-training of modern large language models (LLMs) has been underutilized in general authorship analysis. We introduce an unsupervised framework that uses the log-probabilities of an LLM to measure style transferability between two texts. This framework takes advantage of the extensive CLM pre-training and in-context capabilities of modern LLMs. Our approach avoids explicit supervision with spuriously correlated data. Our method substantially outperforms unsupervised prompting-based baselines at similar model sizes and exceeds contrastively trained models when controlling for topical overlap. Our framework’s performance improves with model size. In the case of authorship verification, we present an additional mechanism that increases test-time computation to improve accuracy; enabling flexible trade-offs between computational cost and task performance.
Keywords Unsupervised $\cdot$ Authorship Attribution $\cdot$ Authorship Verification
1 Introduction
Computational stylometry is the quantitative and computational analysis of writing style, using patterns in vocabulary, syntax, punctuation, and character sequences to identify, verify, or cluster authors based on textual data [daelemans2013ExplanationComputational]. Authorship analysis has been widely applied across diverse domains, with applications spanning both forensic and humanities contexts. In forensics, stylistic and linguistic patterns have been used to analyze terrorist discourse [abbasi2005ApplyingAuthorship], link user identities across online platforms [shu2017UserIdentity], detect plagiarism [stamatatos2011Plagiarismauthorship], and identify coordinated disinformation campaigns [stiff2022Detectingcomputergenerated]. Authorship methods also play a key role in the humanities, supporting tasks such as the attribution of pseudepigraphic texts [koppel2013Automaticallyidentifying] and the analysis of anonymous literary works [stover2016Computationalauthorship].
Authorship analysis is a long-standing challenge in natural language processing, with early approaches that take advantage of hand-crafted stylometric features [stamatatos2016AuthorshipVerification, lagutina2019SurveyStylometric]. The advent of pre-trained transformers [vaswani2017AttentionAll, devlin2019BERTPretraining] led to the use encoder-only models [fabien-etal-2020-bertaa, zangerle2023Overviewmultiauthor]. Yet, such fine-tuned models often struggle in the absence of topical correlations [zangerle2023Overviewmultiauthor].
A different supervised approach in the literature involves fine-tuning individual language models for candidate authors [huang2025Attributingauthorship, barlas2020CrossDomainAuthorship], measuring the probability assigned to texts by the language model. While this approach tries to take advantage of the pre-training of LLMs, it is expensive to apply, and not useful in low resource contexts like authorship verification tasks.
In this work, we focus on non-supervised approaches, given the tendency of supervised training to produce models biased towards semantic similarity. In this category, we include semi-supervised contrastive learning approaches [ai2022WhodunitLearning, huertas-tato2024PARTPretrained, huertas-tato2024Understandingwriting], which still remain limited by the reliance on labeled data. The emergence of large language models (LLMs) has introduced prompt-based unsupervised stylistic analysis [hung2023WhoWrote, huang2024CanLarge], offering an alternative without supervision but facing poor performance (at least at moderate model sizes; see fig.˜5) and challenges in attribution due to context length constraints.
Figure 1: High-level overview of OSST-based methods for authorship attribution and verification. Text highlighted in orange means that we obtain the log probabilities of the corresponding tokens, and {bracketed text} means variable interpolation. The top row defines common operations, while the bottom row defines the application of those operations to authorship verification and attribution.
Proposal.
We propose a novel unsupervised approach to authorship attribution and verification that leverage the extensive CLM pre-training of modern LLMs and their in-context learning capabilities. For a given text $t$, we obtain a version $N(t)$ in neutral style via LLM prompting. Next, we consider the task of predicting the styled text $t$ from its neutral version $N(t)$. This task can be performed by an LLM, and, further, we can provide it with in-context few-shot examples. Given a single in-context one-shot example, the average of the LLM log-probabilities of the target text $t$ (or OSST score) reflects how helpful the style of the one-shot example was. Therefore, we use this metric to measure style similarity between the texts. The OSST score supports zero-shot closed-set attribution, though it requires the selection of a decision boundary for open-set attribution and verification. We provide a more in-depth explanation in fig.˜1 and section˜3.
Contributions.
The main contributions of this work are the following.
  1. 1.
    We propose a novel unsupervised approach to authorship attribution and verification taking advantage of the extensive CLM pre-training of modern LLMs and their in-context learning capabilities.
  2. 2.
    We empirically validated its performance for authorship attribution and verification across several datasets and compared with a contrastively trained and two prompting-based baselines.
  3. 3.
    We studied the performance of our methods across multiple base LLM sizes, and performed ablations on another verification-specific hyperparameter that scales test-time compute.
  4. 4.
    We studied the effect of distribution shifts on the decision boundary used by our methods on authorship verification tasks.
  5. 5.
    We validated the multilingual performance of OSST-based methods in authorship attribution.
The source code is available at pablomiralles22/paper-authorship.
Figure 2: Closed-set authorship attribution results. Average accuracy and rank of correct candidate are reported for each model and dataset.
2 Background
2.1 Authorship tasks
In this work, we focus on two core authorship tasks: verification and attribution. Other authorship tasks, such as style change detection, can be reduced to them, and our method is not suited for authorship profiling.
Authorship verification (AV).
Given a set of texts $T_{a}$ known to be written by the author $a$, the task is to make a binary decision on whether a target text $t$ was written by $a$. A common setting is to have a single text in $T_{a}$, making this the symmetric task of deciding whether two texts are written by the same author.
Authorship attribution (AA).
Given a target text $t$ and a set of candidate authors $A=\{a_{1},a_{2},\dots,a_{n}\}$, the task is to determine the most likely author of $t$. To do this, a set of training texts $T=\{T_{a_{1}},T_{a_{2}},\dots,T_{a_{n}}\}$ is provided, where $T_{a_{i}}$ is a set of texts known to be written by the author $a_{i}$.
This task can be presented in two settings. Closed-set authorship attribution involves the assumption that one of the candidates is the author of the target text. This is not the case in open-set authorship attribution, where the method must be able to detect if none of the candidates authored the target text. This can be seen as a closed-set authorship attribution problem together with an authorship verification one.
2.2 Available data
Authorship datasets are foundational for advancing research in computational authorship analysis. Diverse corpora exist across multiple domains, including news articles [lewis2004RCV1new], opinion pieces [stamatatos2013RobustnessAuthorship], product reviews [seroussi2014AuthorshipAttribution], blogs [schler2006Effectsage], literary fiction [gerlach2020StandardizedProject], and social media platforms such as Reddit [volske2017TLDRMininga] and Twitter [cheng2010Youare]. For robust evaluation, this work focuses on datasets that have been cleaned and standardized for AA and AV tasks.
A significant contribution to dataset quality and diversity comes from the CLEF Initiative’s annual PAN competitions, which set benchmarks for digital text forensics and stylometry. Early AA datasets were introduced in 2011 [argamon_2011_3713246] and 2012 [juola_2012_3713273], drawing from corporate emails and fiction, respectively. The 2018 PAN AA task [kestemont_2018_3737849, kestemont2018overview] established a challenging cross-fandom scenario within fanfiction, where all unknown documents originate from a single fandom, while candidate authors’ documents span non-overlapping fandoms. This design amplifies complexity by introducing domain shift and reducing stylistic overlap, as authors often emulate the source material. The 2019 task [kestemont_2019_3530313, kestemont2019overview] further advanced these challenges through an open-set evaluation.
PAN competitions have similarly contributed standardized AV datasets. The 2020 and 2021 editions [bischoff2020ImportanceSuppressing, bevendorff2020PAN20Authorship] centered on fanfiction, with the latter introducing an out-of-distribution test set featuring unseen authors and novel topics. In 2022 [stamatatos2022PAN22Authorship], the AV task expanded to encompass diverse discourse types, including essays, emails, text messages, and business memos.
Additionally, we leverage AV datasets constructed from style-change detection tasks. These require models to determine whether consecutive paragraphs are authored by the same person. We consider data from the 2022 [zangerle2022PAN22Authorship], 2023 [zangerle2023PAN23MultiAuthor], and 2024 [zangerle2024PAN24MultiAuthor] competitions. The 2022 dataset features StackExchange posts, while the 2023 and 2024 datasets focus on Reddit, specifically curating subsets where input texts revolve around the same topic. This setting limits reliance on semantic cues and accentuates the need for nuanced stylistic detection.
2.3 Authorship methods
Traditional authorship analysis methods relied on handcrafted lexical (e.g., character and word n-grams, word frequency, word-length distribution), syntactic (e.g., POS distributions, syntactic n-grams), and semantic features [stamatatos2016AuthorshipVerification, lagutina2019SurveyStylometric]. These features were typically used in supervised settings with classifiers such as SVMs [stamatatos2016AuthorshipVerification], or with early neural architectures including RNNs [bagnall2016AuthorIdentification] and CNNs [shrestha-etal-2017-convolutional].
With the introduction of transformers [vaswani2017AttentionAll], pre-trained encoder-only models such as BERT [devlin2019BERTPretraining] have been widely adopted for authorship-related tasks [fabien-etal-2020-bertaa, zangerle2023Overviewmultiauthor]. These supervised models primarily capture semantic features, which limits their effectiveness in settings where texts share a common topic [zangerle2023Overviewmultiauthor].
An alternative line of work fine-tunes an individual language model for each candidate author using a corpus of known texts, and attributes authorship by evaluating perplexity under these author-specific models [huang2025Attributingauthorship, barlas2020CrossDomainAuthorship]. However, the need to fine-tune a language model for each author makes these methods expensive and not applicable in low-resource environments like authorship verification tasks.
More recently, there has been growing interest in learning author embeddings, where each text is mapped into a vector representation capturing stylistic properties. These methods often rely on contrastive learning [ai2022WhodunitLearning, huertas-tato2024PARTPretrained, huertas-tato2024Understandingwriting] and assess authorship via similarity in the embedding space. However, such approaches remain dependent on labeled data, often strongly correlated with topical similarity.
The latest trend explores the use of large language models (LLMs) for human-like stylistic analysis based on prompts [hung2023WhoWrote, huang2024CanLarge]. Although promising because of their unsupervised nature, these methods require very large models in attribution settings because of context limitations, and empirical evaluation remains scarce. Our results indicate that, in practice, such approaches yield very limited performance in authorship verification.
2.4 Causal Language Modeling and few-shot learning
Causal Language Modeling (CLM) refers to predicting the next token based on prior tokens in a unidirectional, autoregressive manner, ie, modeling $P(x_{t}\mid x_{<t})$ [bengio2003neuralprobabilistic]. It is the pre-training objective used by GPT-style decoder-only models [radford2019language]. A landmark discovery in large language models (LLMs) is their few‑shot learning ability [wei2022Emergentabilities]. GPT 3 (175B parameters) was shown to perform new tasks with just a few textual examples provided in the prompt, without any explicit gradient updates or fine-tuning, demonstrating competitive performance with fully fine-tuned systems [brown2020LanguageModels]. This emergent capability highlights that massive autoregressive pre-training endows models with flexible, task‑agnostic generalization using only prompt examples.
We hypothesize that models trained using the CLM objective learn about style, as it is a key component in generating coherent continuations of a text. Furthermore, our proposed method relies on LLMs having the ability to transfer the style of one text to another through one-shot demonstrations.
3 One-shot Style Transfer (OSST) with LLMs for authorship attribution and verification
In this section we provide an in-depth explanation of the main components and intuitions behind our OSST-based methods for AA and AV. A high-level overview of all the components is given in fig.˜1.
Obtaining versions of a text in neutral style.
Our methods require a way to generate neutrally styled versions of a given text. To do this, we prompt an instruction-tuned LLM to rewrite the input text in a neutral style, using the template shown in LABEL:lst:neutralize-prompt. More formally, given a text $t$, we consider the random variable $\mathrm{N}_{M,\theta}(t)$ of generating a neutral text with LLM $M$ and sampling parameters $\theta$. For the rest of the work, we omit the subscript and abuse the notation to signify a sample of the random variable.
X-shot style transfer problem set-up.
As introduced earlier, our method relies on the few-shot capabilities of LLMs to transfer the style of one text to another. In particular, for a pair of texts $t_{1}$ and $t_{2}$, consider the following input:
Neutral text: $\mathrm{N}(t_{1})$}
Styled text: $t_{1}$}
Neutral text: $\mathrm{N}(t_{2})$}
Styled text: $t_{2}$}
We task the LLM with predicting the second text’s styled version based on the neutral version and the example from the first text. In this way, the LLM can predict semantic information from the neutral version while obtaining stylistic information from the one-shot example. To measure the degree of transferability from the style of the first text to the second text, we retrieve the log-probabilities from a given LLM:
$\mathrm{OSST}(t_{2}|t_{1})=\log P\left[t_{2}\mid\mathrm{N}(t_{2});\ t_{1},\mathrm{N}(t_{1})\right].$ (1)
Authorship attribution.
Given a target text $t_{target}$ and a set of candidate authors $\mathcal{A}=\{a_{1},a_{2},\dots,a_{n}\}$, the task is to determine the most likely author of $t_{target}$. To do this, a set of training texts $\mathcal{T}=\{T_{a_{1}},T_{a_{2}},\dots,T_{a_{n}}\}$ is provided, where $T_{a_{i}}$ is a set of texts known to be written by the author $a_{i}$.
To perform this task, our method assigns a score to each candidate author as
$\mathrm{score}(a)=\mathop{\mathrm{mean}}_{t\in T_{a}}\ \mathrm{OSST}(t_{target}|t),$ (2)
and the prediction is the candidate with the highest score.
Open-set authorship attribution.
The open-set attribution setting includes the possibility that the target text is not written by any of the candidate authors. Unlike in the closed-set setting, we must select a global score decision boundary for all samples. In other words, the closed-set setting was an apples-to-apples comparison: the target text was always the same. In the open-set one we are subject to the natural differences in probability distribution across target texts (length, topic, genre, etc.).
Fortunately, the OSST score is computed for the same target text across several one-shot examples, one for each candidate text. This allows us to standardize the scores, alleviating the natural discrepancies between texts.
Authorship verification.
Given a set of texts $T_{a}$ known to be written by the author $a$, the task is to make a binary decision on whether a target text $t$ was written by $a$. The simpler version of deciding whether two texts $t_{1}$ and $t_{2}$ were authored by the same person is often considered.
The verification problem presents the same difficulty as the open-set attribution problem. Again, we must compare the score between many different target texts, which requires some form of normalization. We present two options to solve this problem.
  1. 1.
    Zero-shot baseline (nOSST-ZS). We normalize the $\mathrm{OSST}$ score by comparing it with the log-probabilities of the zero-shot problem, without any in-context example. This is shown graphically in fig.˜1, and mathematically as follows:
    $\mathrm{nOSST}_{zs}(t_{2}|t_{1})=\mathrm{OSST}(t_{2}|t_{1})\\ -\log P\left[t_{2}\mid\mathrm{N}(t_{2})\right]$ (3)
    This normalization step presents some problems. Figure˜3 shows the median $\mathrm{nOSST}_{zs}$ score by token position in the PAN23 style change dataset. As we can see, the first few tokens have a much higher median value. This seems natural, as in the zero-shot problem the task is not completely clear during the first tokens. This presents a challenge in comparing texts of different lengths. We also observe differences between models and datasets.
    Figure 3: Median nOOST score (zero-shot baseline) by token position across models and PAN23 Style Change subtasks, computed from the test set.
  2. 2.
    Random sampling baseline (nOSST-RS). To mitigate the previous problems, we can select a fixed set of anchor texts $\mathcal{T}_{anchor}$, and compute the normalized score as
    $\mathrm{nOSST}_{rs}(t_{2}|t_{1})=\mathrm{OSST}(t_{2}|t_{1})-\\ \mathop{\mathrm{mean}}_{t\in T_{anchor}}\mathrm{OSST}(t_{2}|t).$ (4)
    In this case, the difficulty lies in selecting the distribution from which to sample anchor texts.
Since the task is symmetric across texts, we can consider the score
$\mathrm{score}(t_{1},t_{2})=\frac{1}{2}\left(\mathrm{nOSST}_{*}(t_{1}|t_{2})+\mathrm{nOSST}_{*}(t_{2}|t_{1})\right).$ (5)
4 Experimental results
We conducted several experiments with the purpose of evaluating and studying OSST-based methods in attribution and verification.
4.1 Authorship attribution
We evaluated our models in established authorship attribution benchmarks in the fiction domain, including PAN12 [juola_2012_3713273], PAN18 [kestemont_2018_3737849], and PAN19 [kestemont_2019_3530313], and in the corporate email domain with PAN11 [argamon_2011_3713246]. Our main baseline is STAR [huertas-tato2024Understandingwriting], the strongest contrastive learning method reported in prior work. We do not compare against LLM prompting approaches such as LIP [huang2024CanLarge], as these methods require large context windows in AA, resulting in markedly lower performance at moderate model sizes.
Figure˜2 presents the average accuracy and rank of correct candidate for closed-set attribution. In the fiction domain, our models with 4B and 8B parameters outperform the STAR baseline across all datasets. Notably, we observe a consistent increase in performance with larger model sizes, indicating a clear scaling trend. This suggests that further gains may be achievable with continued scaling, although its returns seem to be diminishing. In the PAN11 dataset, all OSST models achieve higher accuracy but exhibit poorer overall ranking of the correct candidate. Interestingly, model performance does not improve with increasing base model size. A likely explanation is the noisier nature of PAN11 compared to other datasets, as it contains erroneous messages, irregular candidate set sizes, and other inconsistencies.
Table˜1 summarizes results in the open-set setting, where all methods operate in a zero-shot regime and produce similarity scores between text pairs. Scores are standardized across candidate texts (for all models) to improve comparability of scores for different target texts. A small held-out subset is used to select the decision threshold. Our models exhibit robust performance in the open-set scenario, with all parameter sizes surpassing STAR. Notably, OSST-based models demonstrate less performance degradation relative to closed-set attribution, as measured by accuracy.
Table 1: Open-set authorship attribution evaluation. Best results are highlighted in bold. BinF1Macro denotes the macro-averaged F1 score for the binary classification of true author presence among candidates, while the average rank of the correct candidate is calculated only for instances where the true author is included. All similarity scores are standardized across candidate texts.
Model Acc BinF1Macro Rank
PAN19-en
[…]
4.2 Authorship verification
For authorship verification (AV), we evaluate on the 2020 and 2021 editions of the PAN AV shared task [bischoff2020ImportanceSuppressing, bevendorff2020PAN20Authorship], which focus on fanfiction. In addition, we construct AV datasets from the PAN style-change detection tasks of 2022–2024 [zangerle2022PAN22Authorship, zangerle2023PAN23MultiAuthor, zangerle2024PAN24MultiAuthor]. The 2022 edition contains StackExchange posts, while the 2023 and 2024 editions comprise Reddit posts and explicitly control for topic, thereby preventing detectors from exploiting semantic information.
As in authorship attribution (AA), we compare against the STAR baseline [huertas-tato2024Understandingwriting]. We include LIP [huang2024CanLarge] and PromptAV [hung2023WhoWrote] with a 24B Mistral LLM as zero-shot baselines. For STAR and OSST-based methods, we select a decision boundary with 2000 examples from the training set.
For nOSST methods, we focus on LLaMA base models, as they consistently outperform Gemma models at comparable scales, and we further include a 24B Mistral model to test larger-capacity base models. In the nOSST-RS setting, we use 10 anchors sampled in-distribution from the training set.
The results on authorship verification datasets are summarized in figs.˜4 and 5. Several key patterns emerge: (i) nOSST consistently surpasses STAR on settings that control for topical correlations (the third tasks in PAN23ST and PAN24ST), even in its zero-shot (ZS) configuration with an 8B model. In contrast, STAR is better in datasets with strong topic biases. (ii) At comparable model scales (up to 24B parameters), nOSST methods substantially outperform LLM-based prompting approaches. (iii) The RS variant of nOSST generally improves performance across multiple datasets, though a few exceptions remain. (iv) Performance exhibits a clear scaling trend: larger nOSST base models tend to achieve higher accuracy, with only minor deviations from monotonic improvement.
Figure 4: F1 score on style change tasks in AV format.
Figure 5: F1 score on authorship verification tasks.
4.3 Effect of the number of anchor samples in nOSST-RS
Figure˜6 presents the mean verification F1 scores on all verification datasets as a function of the number of anchor examples provided to nOSST. The point at zero corresponds to the zero-shot baseline setting, and, for computational efficiency, we re-sample from a fixed pool of 10 examples. Notably, introducing even a single anchor example yields a substantial improvement over the zero-shot baseline with minimal additional cost. While performance gains exhibit a pattern of diminishing returns up to a plateau as the number of anchor examples increases, we observe that the Mistral24B base model shows an upward final trajectory in F1.
Figure 6: Average verification F1 across all datasets by number of anchor examples in nOSST. At 0 we show the performance of zero-shot baseline, and at 1010 a single set of anchor samples was used.
4.4 Decision boundary under distribution shifts in nOSST
Figure˜7 reports cross-dataset performance of nOSST scores for authorship verification, where the decision threshold is tuned on one dataset and evaluated on another. We consider three domains: fanfiction, StackExchange, and Reddit. As expected, thresholds selected in-distribution yield the best results. We observe a reasonable performance degradation in most pairs of datasets. However, this is not the case with the PAN20 dataset. This great difference could be partially explained by the great difference in text lengths (see figs.˜9 and 3), but not fully, as training in PAN22’s first task does generalize to PAN20 in the RS variant.
Figure 7: Cross-dataset macro-F1 of nOSST scores for authorship verification with a LLaMA-8B base model. The training set is used to select the decision boundary.
4.5 Multi-language performance in authorship attribution
Figure 8: Average accuracy and rank of correct candidate by language and method in the PAN18 and PAN19 authorship attribution datasets.
In fig.˜8, we present the average accuracy and rank of correct candidate of the proposed OSST-based method across various languages and backbone models for authorship attribution (AA), using the PAN18 and PAN19 multi-language AA datasets. STAR is included as a baseline for comparison. As anticipated, performance declines for all models outside of English, with the OSST-based method consistently outperforming STAR by a larger margin in non-English settings. This disparity can be attributed to the limited multilingual training of STAR, whereas OSST benefits from pre-training on large-scale cross-lingual corpora. These results show OSST’s superior generalization and robust multilingual capabilities in authorship attribution tasks.
5 Conclusions
We introduced OSST-based methods, a family of unsupervised approaches to authorship attribution and verification that take advantage of the extensive causal language modeling pre-training of modern LLMs. This paradigm enables scalable authorship analysis without relying on labeled data prone to topical biases.
Empirically, OSST-based methods outperform LLM prompting alternatives of comparable scale and surpass contrastively trained baselines when topical correlations are controlled. Our findings also suggest promising scaling behaviors: performance improves with model size and, for the verification random sampling variant (nOSST-RS), with the number of anchor examples, up to a saturation point. These trends indicate a practical trade-off between computational resources and stylistic accuracy.
The present work is limited by model size and by the requirement of in-distribution anchors in the RS variant, which may limit real-world applicability. Future research should explore scaling laws with larger base models and alternative anchor selection mechanisms. One such alternative could be the use of contrastively trained models for retrieval. We believe these directions can advance unsupervised stylistic modeling and bring LLM-based authorship analysis closer to robust, real-world deployment.
Acknowledgments
This work has been supported by the following projects: H2020 TMA-MSCA-DN TUAI project ”Towards an Understanding of Artificial Intelligence via a transparent, open and explainable perspective” (HORIZON-MSCA-2023-DN-01-01, Grant agreement nº: 101168344); by project PCI2022-134990-2 (MARTINI) of the CHISTERA IV Cofund 2021 program; by EMIF managed by the Calouste Gulbenkian Foundation, in the project MuseAI; and by Comunidad Autonoma de Madrid, CIRMA-CAM Project (TEC-2024/COM-404).
Appendix A Model and prompt to obtain a neutrally styled version of a text
We use a google/gemma-3-12b-it [team2025Gemma3] model to obtain neutral versions of a given text. This decision was made for its multilingual capabilities and the inclusion of multilingual texts in some of the datasets used.
The prompt we used is the following:
1Text: [TEXT]
2---
3Paraphrase the previous text using the driest, simplest, and most neutral style possible.
4Remove all stylistic features, emotion, or rhetorical devices. Replace as many words as possible with simpler or more common alternatives.
5Preserve as much of the original meaning as possible. Write in [LANGUAGE]. Output only the rewritten version.
Listing 1: Prompt to obtain a neutrally styled version of a text.
Appendix B Data preprocessing
Processing fanfiction data.
To preprocess the fanfiction data used in the PAN18 and PAN19 authorship attribution tasks, we first remove all lines that do not contain alphabetic characters, thereby eliminating chapter separators that should not inform authorship prediction. Next, we remove line breaks to normalize paragraph formatting across authors, as the frequency of line breaks can vary stylistically. Finally, we truncate text at the paragraph level to a maximum of 384 tokens.
Enron emails:
To preprocess the Enron emails data used in the PAN11 authorship attribution tasks, we remove surrounding whitespace, truncate to 384 tokens, and remove author texts under 20 characters, and select a maximum of 50 texts per author to avoid massive computational costs.
Cleaning style change data.
As some paragraphs in the style change data where purely random symbols, we remove any sample where the paragraph does not contain at least one alphabetic character.
Appendix C Verification dataset lengths
Figure˜9 presents the length distributions (in characters) of texts in different verification datasets after truncation.
Figure 9: Length distributions (in characters) of texts in different verification datasets after truncation.
Appendix D Base models used in OSST
The following models were used as base for the OSST-based methods in this work:
  • google/gemma-3-1b-pt [team2025Gemma3]
  • google/gemma-3-4b-pt [team2025Gemma3]
  • meta-llama/Llama-3.2-1B [grattafiori2024Llama3]
  • meta-llama/Llama-3.2-8B [grattafiori2024Llama3]
  • mistralai/Mistral-Small-24B-Base-2501
Appendix E Hardware and software
All experiments were run on 1-3 NVIDIA GeForce RTX 3090 GPUs with 24GB of memory. We used PyTorch 2.6.0 with CUDA 12.4. All source code is available at pablomiralles22/paper-authorship, as well as the versions of the libraries as an uv environment.
Appendix F Do we need neutral examples?
A natural question is whether conditioning on a neutral rewrite is actually required. To assess this, we compare our approach against a baseline (NN) that directly continues the input text without access to a neutral variant. As shown in fig.˜10, omitting the neutral version has little impact (and can even be advantageous) on datasets where topical cues are strongly aligned. However, in settings with minimal topical correlation, performance drops sharply when the task is not cast as neutralization. This is unsurprising: generating a continuation of one text conditioned on another is substantially more difficult when their topics diverge, especially in the initial tokens, which makes distinguishing related from unrelated pairs significantly easier.
Figure 10: Comparison between using (nOSST) and not using (nOSST-NN) a neutral version to compute log-probabilities.
Appendix G To what extent does the neutralizer model influence performance?
In this work we have used a model of 12B parameters to create neutral versions of the given texts. This choice greatly influences the amount of resources needed to run our model overall. Is this an important choice? In fig.˜11 we show a comparison between using a 4B and a 12B model in PAN22-24ST tasks. As we can see, this downgrade does affect performance slightly, but not to a great extent.
Figure 11: Comparison between using a google/gemma-3-4b-it google/gemma-3-12b-it [team2025Gemma3] model to create neutral versions of texts in PAN22ST, PAN23ST and PAN24ST.